The demand equation for a product is given by
P = 30 - 0.1q^2 .
(a)At what point price is demand inelastic?
(b)at what price is demand unitary elastic
1
Expert's answer
2019-01-03T11:38:16-0500

(a) At what point price is demand inelastic?

Answer:

P for q = 15 is;

P=30-0.1*〖15〗^2⇨$7.5

E_p=-1*7.5/15⇨-0.5

As demand is inelastic, when price is $7.5 and quantity demanded is 15.


(b) At what price is demand unitary elastic

Answer:

P=30-0.1q^2

MR=30-0.3q^2

MR = 0, so q is:

30-0.3q^2=0

q^2 =30/0.3=100

q = 10

Then, demand unitary elastic when price is:

P=30-0.1〖*10〗^2

P = $20